Suppose that Maria has 130 coins consisting of pennies, nickels, and dimes. The number of nickels she has is 9 less than twice the number of pennies; the number of dimes she has is 23 less than three times the number of pennies. How many coins of each kind does she have?
Answer by mananth(16949)

Suppose that Maria has 130 coins consisting of pennies, nickels, and dimes. The number of nickels she has is 9 less than twice the number of pennies; the number of dimes she has is 23 less than three times the number of pennies. How many coins of each kind does she have?
Pennies = x numbers
Nickels = 2x-9
Dimes = 3x-23

Add them up
x+2x-9+3x-23=130
6x-32=130
6x=130+32
6x=162
/6
x=27 ----------- pennies
substitute x to get the number of nickels and dimes